OpenStudy (anonymous):

x/5 > or equal to -9

OpenStudy (anonymous):

multiply by 5

OpenStudy (anonymous):

\[\frac{ x }{ 5 } \ge -9\]

OpenStudy (anonymous):

i did but I don't know what way the sign goes

OpenStudy (anonymous):

So What You Want To Do Is Isolate x So, Multiply 5 By Both Sides. \[5*\frac{ x }{ 5 }\ge -9*5\]

OpenStudy (anonymous):

Can You Do That?